Asian Black Latino White
Sample Mean 83.70 73.50 80.60 79.00
Sample Variance 26.30 27.50 28.50 24.80
Sample Size 20 20 20 20

Part A.

According to The Boston Globe (August 16, 2010), Asian residents in Boston have the highest average life expectancy of any racial or ethnic group. The report shows, in particular, that Asians live a decade longer than black residents. Suppose the sample results below are indicative of the overall results. Calculate the F statistic.

  1. 79.20
  2. 13.67
  3. 2.72
  4. 26.70
  5. None of the above

Part B.

Suppose the p-value for the above study = 0.00. Which of the following conclusions with respect to the study is correct?

  1. At the 5% level, do not reject the null hypothesis. Average life expectancy differs across the four racial groups
  2. At the 1% level, reject the null hypothesis. Average life expectancy is the same across the four racial groups
  3. At the 1% level, reject the null hypothesis. Average life expectancy differs across the four racial groups
  4. None of the above
Answer #1

Solution:

Part A.

According to The Boston Globe (August 16, 2010), Asian residents in Boston have the highest average life expectancy of any racial or ethnic group. The report shows, in particular, that Asians live a decade longer than black residents. Suppose the sample results below are indicative of the overall results. Calculate the F statistic.

Answer: d. None of the above.

Part B.

Suppose the p-value for the above study = 0.00. Which of the following conclusions with respect to the study is correct?

Answer: c. At the 1% level, reject the null hypothesis. Average life expectancy diers across the four racial groups

Explanation:

We have to find the ANOVA table to find the F test statistic:

Correction factor:

T, ni 〉-X, 20 x 83.70+20 x 73.50+20 x 80.60+20 x 79 6336

G63365018112 80

small RSS=sum n_{i}left ( s_{i}^2+ar{X_{i}}^2 ight )

20(26.3- 83.7) 20 (27.5-73.52) 20(28.5-80.620(24.8- 792)

505048

Therefore, the total sum of squares is:

TSS- RSS-CF 505048-501811.2-3236.8

The Grand mean is given below:

83.7 73.580.6+79 79.2

Now the sum of squares between groups is given below:

small TrSS=sum n_{i}(ar{x_{i}}-ar{ar{x}})^2

small =20(83.7-79.2)^2+20(73.5-79.2)^2+20(80.6-79.2)^2+20(79-79.2)^2

1094.8

And now Error sum of squares is given below:

small ESS=TSS-TrSS

  3236.8-1094.8

  2142

Now the degrees of freedom:

The degrees of freedom for the total is small N-1=80-1=79

The degrees of freedom for the Between groups is small k-1=4-1=3

The degrees of freedom for the within groups is small (N-1)-(k-1)=79-3=76

Mean sum of squares:

small MSTr=rac{TrSS}{df}=rac{1094.8}{3}=364.9333

TrSS 2142 MSE= = 28.1842 df

Now the F statistic:

small F-statistic = rac{MSTr}{MSE}=rac{364.9333}{28.1842}=12.948

Therefore, none of the options is correct